# Weighted Birkhoff averages: Deterministic and probabilistic perspectives

### v2

@@ 4 Proof of the probabilistic results
$\sum\limits_{n=1}^{N}{\theta_{n}^{2}}=\sum\limits_{n=1}^{N}{w_{p,q}\left({n/N}\right)/{A^{p,q}_{N}}}=\sum\limits_{n=1}^{N}{w_{p,q}\left({n/N}\right)}\Big{/}\sum\limits_{n=0}^{N-1}{w_{p,q}\left({n/N}\right)}=1,\;\;\forall N\geqslant 3.$ (4.1)
@@
${A^{p,q}_{N}}=\sum\limits_{n=0}^{N-1}{{w_{p,q}}\left({n/N}\right)}=N\sum\limits_{n=0}^{N-1}{{N^{-1}}{w_{p,q}}\left({n/N}\right)}\sim N\int_{0}^{1}{{w_{p,q}}\left(x\right)dx}=N,\;\;N\to+\infty.$ (4.2)
@@
5 Appendix
5.1 Orthogonality of other column vectors in $\mathscr{V}$
Proof.
We first show the orthogonality of $\mathscr{V}_{2l-1}$} and $\mathscr{V}_{2j-1}$}, where $l\neq j$ and $l+j\leqslant m+\left({m-1}\right)=2m-1<2m$. Then ${\alpha_{1}},{\alpha_{2}}\in\left({0,\pi}\right)$ (recall that ${\alpha_{1}}=:\pi\left({l+j}\right)/m$ and ${\alpha_{2}}=:\pi\left({l-j}\right)/m$), and
Similarly, for ${\mathscr{V}_{2l}}\cdot{\mathscr{V}_{2j}}$ with $l\neq j$ and $l+j\leqslant 2m-2<2m$, one can verify that
Finally, (3.9) provides the orthogonality of $\mathscr{V}_{0}$ and other column vectors. ∎
@@ Acknowledgements
The first author sincerely thanks Prof. James D. Meiss (University of Colorado) for his generous sharing of the latest developments in weighted Birkhoff averages through insightful discussions. The first author also sincerely thanks Prof. Valery V. Ryzhikov (Moscow State University) for his discussions on the weighted rates of general dynamical systems. The first author is also very grateful to Profs. Ai Hua Fan (Université de Picardie) and Konstantin M. Khanin (University of Toronto) for the discussions, which have helped to extend some of the ideas from the initial version of this paper. These extensions will be discussed in a separate work. The first author sincerely thanks Prof. Bo’az B. Klartag (Weizmann Institute of Science) for providing valuable information on the advances in weighted central limit theorems. The first author sincerely thanks Profs. Aleksandr G. Kachurovskiĭ (Sobolev Institute of Mathematics), Ivan V. Podvigin (Sobolev Institute of Mathematics) and Naci Saldi (Bilkent University) for their invaluable assistance and discussions. This work was supported in part by the National Natural Science Foundation of China (Grant Nos. 12071175 and 12471183).

### v3

@@ 4 Proof of the probabilistic results
$\sum\limits_{n=1}^{N}{\theta_{n}^{2}}=\sum\limits_{n=1}^{N}{w_{p,q}\left({n/N}\right)/{A^{p,q}_{N}}}=\sum\limits_{n=1}^{N}{w_{p,q}\left({n/N}\right)}\Big/\sum\limits_{n=0}^{N-1}{w_{p,q}\left({n/N}\right)}=1,\quad\forall N\geqslant 3.$ (4.1)
@@
${A^{p,q}_{N}}=\sum\limits_{n=0}^{N-1}{{w_{p,q}}\left({n/N}\right)}=N\sum\limits_{n=0}^{N-1}{{N^{-1}}{w_{p,q}}\left({n/N}\right)}\sim N\int_{0}^{1}{{w_{p,q}}\left(x\right)dx}=N,\quad N\to+\infty.$ (4.2)
@@
Declarations
Conflict of interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data availability statements
Data will be made available on reasonable request.
@@ Acknowledgements
The authors would like to sincerely thank the anonymous referees for their valuable suggestions and comments, which significantly improved the paper. Z. Tong sincerely thanks Prof. James D. Meiss (University of Colorado) for his generous sharing of the latest developments in weighted Birkhoff averages through insightful discussions. Z. Tong also sincerely thanks Prof. Valery V. Ryzhikov (Moscow State University) for his discussions on the weighted rates of general dynamical systems. Z. Tong sincerely thanks Prof. Alessandra Celletti (Università Roma Tor Vergata) for her 2024 talk related to weighted Birkhoff averages. Z. Tong is also very grateful to Profs. Ai Hua Fan (Université de Picardie) and Konstantin M. Khanin (University of Toronto) for the discussions, which have helped to extend some of the ideas from the initial version of this paper. These extensions will be discussed in a separate work. Z. Tong sincerely thanks Prof. Bo’az B. Klartag (Weizmann Institute of Science) for providing valuable information on the advances in weighted central limit theorems. Z. Tong sincerely thanks Profs. Aleksandr G. Kachurovskiĭ (Sobolev Institute of Mathematics), Michael Lin (Ben-Gurion University), Ivan V. Podvigin (Sobolev Institute of Mathematics) and Naci Saldi (Bilkent University) for their invaluable assistance and discussions. Z. Tong was supported by the China Postdoctoral Science Foundation (Grant No. 2025M783102). Y. Li was supported in part by the National Natural Science Foundation of China (Grant Nos. 12071175, 12471183 and 12531009).